OpenStudy (anonymous):

Given f(x) = 1/(x ) what is f-1(x) ?

Parth (parthkohli):

\[y = \dfrac{1}{x}\]Replace \(y\) with \(x\) and solve for \(y\).\[x = \dfrac{1}{y} \Rightarrow xy = 1 \Rightarrow y = \dfrac{1}{x}\]